Hi
Anna had some cupcakes. She had 72 more chocolate than red velvet, 36 fewer blueberry than red velvet. After selling 1/6 of the chocolate 2/3 of the red velvet and 7/9 of the blueberry she had 427 cupcakes left. How many chocolate cupcakes did she sell. Thanks

Answer by Boreal(15235)

RV=x
C=x+72
B=x-36
(5/6)(x+72)+(1/3)x+(2/9)(x-36)=427
(50/36)x+52=427
(50/36)x=375
x=270
RV=270
C=342
B=234
sold 1/6 of the chocolate which was 57 of them.
-check
she had 846 cupcakes
sold 57 chocolate, 180 red velvet, and 182 blueberry, which is 419. Subtract that from 846 and she had 427.